Consider the following algorithm (known as Horner'
Answer #1

a) f(x)=4x+8x+x+2 for x=3

= 2+x(4+8+1)

= 2+3(13)

= 2+39 =41

b) To convert between different positional of numeral systems we will use this method. In this method, case x is the base of the number system, and the ai coefficients are the digits of the x representation for a given number. And can also be used if x is a matrix, where the gain in computational efficiency is large. Actually, when x is a matrix, further acceleration is possible where we look deep into the structure of matrix multiplication, where only {\displaystyle {\sqrt {n}}}instead of n multiplies are needed (at the expense of requiring more storage)

c) O(n)= as we go with polynomial equation